PRESSURE BROADENING IN THE ^{13}C^{16}O 2–0 BAND

\underline{S. VOIGT}, S. DREHER, J. ORPHAL, AND J. P. BURROWS, Institute of Environmental Physics, University of Bremen, PO Box 330440, D-28334 Bremen, Germany.

N_{2} induced pressure broadening in the ^{13}C^{16}O 2–0 band around 4167 cm^{−1} was investigated using Fourier Transform Spectroscopy. The spectra were recorded with a sample of 99 % isotopically pure ^{13}CO in a White-type absorption cell at a resolution of 0.005 cm^{−1}. Voigt profiles convolved with the FTS apparatus function were fitted to the observed lineshapes^{a}, and Lorentzian HWHM were determined as function of N_{2} pressure. Pressure broadening coefficients for m between –33 and +34 were obtained with uncertainties of 5.8 %.^{b}. The results are compared to N_{2} broadening coefficients obtained from simultaneous measurements in the 2–0 band of ^{12}C^{16}O. While the latter values agree well with those published earlier, those of^{13}C^{16}O are systematically lower by 5–7 % compared to ^{12}C^{16}O. This indicates that in spectroscopic databases the pressure broadening of ^{13}C^{16}O is significantly overestimated. This is most important since recent work^{c} shows that the ^{13}CO/^{12}CO mixing ratio is an indicator of stratospheric O_{3} depletion chemistry.

aK. Chance, Smithsonian Center for Astrophysics, Cambridge, USA, personal communication (1995).

bS. Voigt et al., J. Mol. Spectrosc. 180, 359 (1996).

cC. A. Brenninkmeijer et al., Geophys. Res. Lett. 23 (16), 2125 (1996).